Demography & International Capital Flows

David Backus, Thomas Cooley, and Espen Henriksen NYU | Economics Honors Society April 4, 2012

This version: May 10, 2012

Backus, Cooley, and Henriksen (NYU)  International Capital Flows  1 / 29

Capital flows: current accounts

Largest deficits Amount (b$) Largest surpluses Amount (b$)
United States  –470 China  306
Italy  –72 Japan  195
Spain  –63 Germany  176
United Kingdom  –56 Mideast oil  150
France  –53 Switzerland  75
India  –49 Russia  71
Turkey  –49 Netherlands  56
Canada  –49 Norway  53
Brazil  –47 Taiwan  41
Greece  –32 Sweden  30
Australia  –32 Korea  28
Portugal  –23 Malaysia  28
Total deficits  –1183 Total surpluses  1466

Source: IMF, WEO, April 2011 version, data for 2010

Capital inflows: sign of success?

Larry Summers, IMF, October 3, 2004:

There is a standard set of things that finance ministers of countries with significant current account deficits say. Perhaps the sharpest formulation is: “We live in a country that capital is trying to get into. Would you rather live in a country that capital is trying to get out of?”

Capital inflows: portent of peril?

Daniel Gross, New York Times, May 8, 2005

[US] imbalances are eerily reminiscent of recent economic crises. Could we see a perfect storm [for the US economy]? If so, what would it look like?

⋆ Nouriel Roubini estimates that long-term interest rates in the US could rise sharply and the dollar fall.

⋆ Jeffrey Frankel adds: “some of us have been warning of this hard-landing scenario for more than 20 years.”

Øystein Olsen, Norges Bank, March 2011

Global trade imbalances have been reduced somewhat over the past two years, but there is a considerable risk that they will persist. They must be corrected.

Where we’re headed

Suggestion: global find and replace

▶ “global imbalances” → “international capital flows”

Our question: where do capital flows come from?

▶ Relate: capital flows, demography, and China

Plan of attack

▶ Evidence ▶ Model ▶ Questions about China

Evidence

Evidence: current accounts

Evidence

Evidence: net foreign assets

Evidence

Evidence: current account spectra

Evidence

Evidence: summary for capital flows and stocks

Capital flow variation is low-frequency [so much for HP filtering] Ditto net foreign assets Needed: an explanation for low-frequency movements

Evidence

Evidence: demography

Evidence

Evidence: summary

Capital flows are persistent

▶ Ditto net foreign assets

Demography inherently persistent

▶ And there are sharp differences across countries

Worth exploring a connection? [Apparently yes: Attanasio-Kitao-Violante, Bloom-Canning, Boersch-Supan-Winter, Brooks, Domeij-Floden, Feroli, Ferrero, Henriksen, Krueger-Ludwig, and others had the same idea]

Model

Model: mechanics

One-good deterministic world Countries: G7 (for now) Each has OG structure Power utility, fixed labor supply [+ age-specific adjustments] Identical production functions, country-specific productivity Fixed retirement age: 65 No bequests: perfect insurance for mortality risk No frictions either within or across countries

Model

Model: how it works

NFA is difference between net worth and capital stock

▶ Think of each as a function of age ▶ Aggregates are sums across age groups

Net worth follows from life-cycle saving

▶ Net worth varies by age ▶ Aggregate net worth depends on age distribution

Capital allocated to equate marginal products

▶ Connected to age distribution via effective labor input ▶ Equates K/Y across countries

Model

Model: questions

Individual saving rates differ across countries

▶ Not clear why

▶ Should we add “wedges” to mimic?

Capital-output ratios differ across countries

▶ Ditto

Market value ≠ capital stock

▶ Ditto

Model

Model: parameter values

Time interval  one year
Capital share (α)  1/3
Risk aversion (σ)  4
Discount factor (β)  1/1.025
Depreciation rate (δ)  2.5%
Fertility and mortality  UN data
Cohort sizes  UN data

Backus, Cooley, and Henriksen (NYU)  International Capital Flows  14 / 29

Model

Model: net foreign assets

Open questions

What about China?

Investment

▶ Capital-output ratio seems fine

Saving

▶ We can explain a high saving rate ▶ But not as high as we see ▶ What’s missing?

Open questions

Why are China’s saving and investment so high?

Open questions

Does China have too much capital?

Open questions

Does China have too much capital?

No! Capital-output ratio not unusual Why not?

▶ PWT implies part of investment is high prices ▶ Growth central to steady state capital-output ratio

K/Y = \frac{I /Y}{g + δ}

[When you grow this fast, you need a lot of investment]

Open questions

Does China have too little capital?

Stories — and a look around — suggest lots of capital projects Also shortage of funds for private firms Can both be true?

Open questions

Why is saving so high?

Demography gets us only part way Precautionary saving? High “business saving”?

Open questions

Is convertibility an issue?

Category  Assets (tr$) Liabilities (tr$)
Direct investment  0.230  0.997
Portfolio investment  0.243  0.190
Other  0.535  0.451
Official reserves  2.453  —
Total  3.460  1.638
Net foreign assets  1.822

Source: IMF, IFS, May 2011 version, data for 2009

Open questions

Is China like Norway?

Norwegian Petroleum Fund (“Pension Fund Global”)

▶ Oil royalties not income to households ▶ Invested abroad by government

Story about China

▶ Business saving primarily SOEs ▶ Not income to households ▶ Invested abroad by government ▶ Like Norway?

Final thoughts

Last thoughts

Capital flows

▶ A fact of life ▶ Persistent ▶ A role for demography?

Open questions

▶ Why are saving rates so different across countries? ▶ Why is China’s so high?

Extra slides

China: GDP

Extra slides

China: GDP per person

Extra slides

China: growth

Extra slides

US: household net worth

Extra slides

US: personal consumption